At three corners of a rectangle (length 14.8 cm, height 7.4 cm), the following charges are located: q1 (upper left corner), q2 (lower right corner), and q (lower left corner). The net electric field at the (empty) upper right corner is zero. Find the magnitudes of (a) q1 and (b) q2 if q = 8.5 C.
Answer #1

distance of q1 and q2 from that point is r1=r2=R=14.8 cm

distance of q from that point is r=(14.82+7.42)1/2

if the eclectric filed at that point is zero, it means

resultant field of q1 and q2 is equal and opposite to q

it means both q1 and q2 are of -ve, and of equal magnitudes

so let q1=q2=Q

E=E1E2sin90

=E1E2

E=(kQ/R2)2

Eq=kq/r2

now

E=Eq

we know all values,we can find Q